# How Effective Are Time-Series Models for Rainfall Nowcasting? A Comprehensive Benchmark for Rainfall Nowcasting Incorporating PWV Data

### v2

@@ Abstract
Rainfall nowcasting, which aims to predict precipitation within the next 0 to 3 hours, is critical for disaster mitigation and real-time response planning. However, most time series forecasting benchmarks in meteorology are evaluated on variables with strong periodicity, such as temperature and humidity, which fail to reflect model capabilities in more complex and practically meteorology scenarios like rainfall nowcasting. To address this gap, we propose RainfallBench, a benchmark designed for rainfall nowcasting, a highly challenging and practically relevant task characterized by zero inflation, temporal decay, and non-stationarity, focused on predicting precipitation within the next 0 to 3 hours. The dataset is derived from five years of meteorological observations, recorded at 15-minute intervals across six essential variables, and collected from more than 12,000 GNSS stations globally. In particular, it incorporates precipitable water vapor (PWV), a crucial indicator of rainfall that is absent in other datasets. We further design specialized evaluation strategies to assess model performance on key meteorological challenges, such as multi-scale prediction and extreme rainfall events, and evaluate over 20 state-of-the-art models across six major architectures on RainfallBench. Additionally, to address the zero-inflation and temporal decay issues overlooked by existing models, we introduce Bi-Focus Precipitation Forecaster (BFPF), a plug-and-play module that incorporates domain-specific priors to enhance rainfall time series forecasting. Statistical analysis and ablation studies validate the comprehensiveness of our dataset as well as the superiority of our methodology. Code and datasets are available at https://anonymous.4open.science/r/RainfallBench-A710.
Figure 1: The Overall Architecture of RainfallBench.
Introduction
@@
Rainfall nowcasting, which focuses on predicting precipitation within the next 0 to 3 hours (Zhang et al. 2023; Nearing et al. 2024), plays a crucial role in disaster mitigation, flood prevention, and real-time decision-making in weather-sensitive sectors. However, current time series forecasting models in meteorology are often evaluated on variables that exhibit strong periodic patterns, such as temperature and humidity. While these benchmarks facilitate model development and comparison, they often fall short in capturing the complexity and uncertainty inherent in real-world meteorology scenarios, and do not adequately assess model performance on rainfall prediction—one of the most critical atmospheric variables. This gap raises concerns about the practical applicability and robustness of existing models.
@@
To bridge this discrepancy, we introduce RainfallBench, a benchmark tailored for rainfall nowcasting — a task characterized by zero inflation, temporal decay, and non-stationarity arising from complex atmospheric dynamics. These properties pose substantial challenges to time-series models, making rainfall nowcasting a more realistic and demanding benchmark for evaluating their effectiveness in practical scenarios.
@@
In recent years, rainfall forecasting has spurred intense activity from the deep learning community. On one hand, most precipitation nowcasting methods rely on weather radar imagery (Franch et al. 2020; Tang et al. 2023; Rivero, Patiño, and Pucheta 2015; An et al. 2025), which is effective but constrained by high costs, limited coverage, and inconsistent continuity. On the other hand, existing benchmarks for time-series forecasting models in in meteorology mainly target longer-term multivariate climate variable forecasting (Wu et al. 2021; Mouatadid et al. 2023) and do not address rainfall nowcasting needs.
@@
In particular, when using the commonly adopted Weather dataset  [^1] https://www.bgc-jena.mpg.de/wetter/ for time series forecasting, most models adopt a multivariate prediction setting, and even in univariate settings, the target variable is typically the last column—CO2 concentration of ambient air. As a result, the evaluation metrics derived from this dataset do not adequately reflect the capability of time series models in the context of rainfall forecasting. Moreover, effective rainfall nowcasting depends on variables strongly correlated with precipitation (e.g., PWV) over the nowcasting horizon, which are often missing from other datasets, making model development and evaluation challenging. To the best of our knowledge, RainfallBench is the first benchmark dedicated to rainfall nowcasting solely based on historical numerical meteorological records, explicitly incorporating PWV.
@@
It consists of data collected from over 12,000 GNSS stations globally between 2018 and 2022, covering six key meteorological variables. All measurements are sampled at 15-minute intervals, enabling fine-grained temporal modeling.
@@
Specifically, RainfallBench offers several distinctive characteristics that set it apart from existing rainfall forecasting datasets: i) Integration of GNSS-Derived Atmospheric Water Vapor: The dataset includes precipitable water vapor (PWV) derived from GNSS observations, which reflects atmospheric moisture and correlates strongly with precipitation onset, making it a key and timely indicator for short-term rainfall nowcasting (Yin et al. 2024). ii) High-Resolution Temporal Sampling: All variables are recorded at 15-minute intervals, enabling the capture of rapid atmospheric dynamics. This high-frequency sampling improves the suitability of the dataset for rainfall nowcasting within a 0 to 3 hours horizon. iv) Derived from Latest Real-World Scenarios: Our collected dataset comprises records from 2018-2022, obtained from professional meteorological observation stations, ensuring its strong practical utility.
@@
To ensure a professional evaluation, we propose a more holistic evaluation framework, which assesses models across three key dimensions: i) Multi-Time Scale Prediction Evaluation: It measures a model’s ability to predict rainfall at various temporal granularities. ii) Extreme Rainfall Event Evaluation: It focuses on a model’s performance in forecasting sudden, high-intensity rainfall events. iii) Accuracy and Reliability of Time Series Prediction: It comprehensively considers both the precision and stability of the model’s predictions. Based on this framework, we benchmark over 20 state-of-the-art deep learning models spanning various architectural families.
@@
Through a comprehensive evaluation, we identify that existing models often overlook the zero-inflation and temporal decay characteristics, compared to the widely acknowledged non-stationarity of time series data (Liu et al. 2022, 2023b, 2025b, 2025a). To address these limitations, we design the BFPF to reinforce its sensitivity to rainfall patterns and recent temporal information. Experimental results validate the effectiveness of our approach, offering a new perspective for adapting time series models to rainfall nowcasting.
@@
  • Professional Dataset for Rainfall Nowcasting: Our benchmark is constructed from data collected at 15-minute intervals between 2018 and 2022 from over 12,000 GNSS stations globally. It covers six key variables, including PWV, and is specifically curated to support rainfall nowcasting. The dataset will be continuously updated.
@@
  • Rainfall-Centric Evaluation Strategy: We design a tailored evaluation strategy from a meteorological perspective, focusing on multi-scale forecasting and extreme rainfall events.
@@
  • Novel plug-and-play Module for Rainfall Nowcasting: We introduce the Bi-Focus Precipitation Forecaster, a plug-and-play module that explicitly addresses zero inflation and temporal decay in rainfall data, achieving state-of-the-art performance in extreme rainfall forecasting.
To facilitate subsequent data analysis and experiments, we select one representative dataset, hereafter referred to as JFNG, as the primary example throughout this study. Data statistics for other stations are provided in the supplementary material.
Related Works
Benchmarks for Time-Series Forecasting
@@
FinTSB (Hu et al. 2025a) emphasizes diversity, standardization, and real-world relevance in financial forecasting. Physiome-ODE (Klötergens et al. 2025) introduces irregularly sampled ODE-based biological datasets for IMTS evaluation. Cherry-Picking (Roque et al. 2025) warns against dataset bias and calls for more representative evaluations. TSFM-Bench and GIFT-Eval Aksu et al. (2024) assess foundation models in zero-, few-, and full-shot regimes. TFB (Qiu et al. 2024) and TSPP (Bączek et al. 2023) propose unified pipelines to ensure fair and reproducible forecasting. LargeST (Liu et al. 2023a) offers a long-term, large-scale traffic dataset with rich metadata to evaluate deep models in realistic settings.
Time-Series Models in Meteorology
PostRainBench (Tang et al. 2023) targets NWP post-processing using deep models. RainBench (de Witt et al. 2021) provides a large-scale, multi-modal benchmark using SimSat, ERA5, and IMERG for global precipitation forecasting. Rodriguez Rivero et al. (Rivero, Patiño, and Pucheta 2015) propose a neural method for short-term rainfall prediction under missing data, leveraging time series energy and the Hurst parameter. Ana et al. (An et al. 2025) provides a comprehensive review of deep learning-based precipitation forecasting methods that utilize multi-source observational data, such as radar reflectivity and satellite imagery. Compared to time-series models with observational data, radar imagery-based methods often entail limited coverage and complex preprocessing, hindering practical deployment.
Unlike benchmarks that rely on multi-modal input, RainfallBench is tailored for rainfall nowcasting, evaluating models solely on historical meteorological time series enriched with key signals such as PWV.
RainfallBench
@@
RainfallBench is structured into three main components: the data layer, the model layer, and the evaluation layer. The overall framework is illustrated in Figure 1.
Problem Definition
@@
We formulate rainfall nowcasting as a multivariate-to-univariate time series prediction task. Given a sequence of historical observations comprising both meteorological variables and past rainfall values, the goal is to predict future rainfall over a fixed horizon.
@@
where $\mathbf{x}_{t}$ includes meteorological factors (e.g., temperature, humidity, wind) and the rainfall measurement at time $t$, and $D$ denotes the number of input variables.
@@
Datasets
Data Overview
Figure 2: Variable Distributions

### v3

# How Effective Are Time-Series Models for Precipitation Nowcasting? A Comprehensive Benchmark for GNSS-based Precipitation Nowcasting

@@ Abstract
Precipitation Nowcasting, which aims to predict precipitation within the next 0 to 6 hours, is critical for disaster mitigation and real-time response planning. However, most time series forecasting benchmarks in meteorology are evaluated on variables with strong periodicity, such as temperature and humidity, which fail to reflect model capabilities in more complex and practically meteorology scenarios like precipitation nowcasting. To address this gap, we propose RainfallBench, a benchmark designed for precipitation nowcasting, a highly challenging and practically relevant task characterized by zero inflation, temporal decay, and non-stationarity, focusing on predicting precipitation within the next 0 to 6 hours. The dataset is derived from five years of meteorological observations, recorded at hourly intervals across six essential variables, and collected from more than 140 Global Navigation Satellite System (GNSS) stations globally. In particular, it incorporates precipitable water vapor (PWV), a crucial indicator of rainfall that is absent in other datasets. We further design specialized evaluation protocols to assess model performance on key meteorological challenges, including multi-scale prediction, multi-resolution forecasting, and extreme rainfall events, benchmarking 17 state-of-the-art models across six major architectures on RainfallBench. Additionally, to address the zero-inflation and temporal decay issues overlooked by existing models, we introduce Bi-Focus Precipitation Forecaster (BFPF), a plug-and-play module that incorporates domain-specific priors to enhance rainfall time series forecasting. Statistical analysis and ablation studies validate the comprehensiveness of our dataset as well as the superiority of our methodology.
Figure 1: Overview of the RainfallBench framework. The benchmark is organized into three main components: the data layer, the model layer, and the evaluation layer. The data layer integrates three sources: GNSS-PWV, ERA5, and GPM. The model layer includes 17 models across six major deep learning architectures, while the evaluation layer encompasses multi-scale prediction, multi-resolution forecasting, and extreme rainfall assessment.
I Introduction
@@
Precipitation nowcasting, which focuses on predicting precipitation within the next 0 to 6 hours [1, 2], plays a crucial role in disaster mitigation, flood prevention, and real-time decision-making in weather-sensitive sectors. However, current time series forecasting models in meteorology are often evaluated on variables that exhibit strong periodic patterns, such as temperature and humidity. While these benchmarks facilitate model development and comparison, they often fall short in capturing the complexity and uncertainty inherent in real-world meteorology scenarios, and do not adequately assess model performance on rainfall prediction—one of the most critical atmospheric variables. This gap raises concerns about the practical applicability and robustness of existing models.
@@
To bridge this discrepancy, we introduce RainfallBench, a benchmark tailored for precipitation nowcasting — a task characterized by zero inflation, temporal decay, and non-stationarity arising from complex atmospheric dynamics. These properties pose substantial challenges to time-series models, making precipitation nowcasting a more realistic and demanding benchmark for evaluating their effectiveness in practical scenarios.
@@
In recent years, rainfall forecasting has spurred intense activity from the deep learning community. On one hand, most precipitation nowcasting methods rely on weather radar imagery [3, 4, 5, 6], which is effective but constrained by high costs, limited coverage, and inconsistent continuity. On the other hand, existing benchmarks for time-series forecasting models in meteorology mainly target longer-term multivariate climate variable forecasting [7, 8] and do not address precipitation nowcasting needs.
@@
In particular, when using the commonly adopted Weather dataset  [^1] https://www.bgc-jena.mpg.de/wetter/ for time series forecasting, most models adopt a multivariate prediction setting, and even in univariate settings, the target variable is typically the last column—CO2 concentration of ambient air. As a result, the evaluation metrics derived from this dataset do not adequately reflect the capability of time series models in the context of rainfall forecasting. Moreover, effective precipitation nowcasting depends on variables strongly correlated with precipitation (e.g., PWV) over the nowcasting horizon, which are often missing from other datasets, making model development and evaluation challenging. To the best of our knowledge, RainfallBench is the first benchmark dedicated to precipitation nowcasting solely based on historical numerical meteorological records, explicitly incorporating PWV.
@@
It consists of data collected from over 140 GNSS stations globally between 2018 and 2024, covering six key meteorological variables. All measurements are sampled at hourly intervals, enabling fine-grained temporal modeling.
@@
Specifically, RainfallBench offers several distinctive characteristics that set it apart from existing rainfall forecasting datasets: i) Integration of GNSS-Derived Atmospheric Water Vapor: The dataset includes precipitable water vapor (PWV) derived from GNSS observations, which reflects atmospheric moisture and correlates strongly with precipitation onset, making it a key and timely indicator for precipitation nowcasting [9]. ii) High-Resolution Temporal Sampling: All variables are recorded at hourly intervals, enabling the capture of rapid atmospheric dynamics. This high-frequency sampling improves the suitability of the dataset for precipitation nowcasting within a 0 to 6 hours horizon. iv) Derived from Latest Real-World Scenarios: Our collected dataset comprises records from 2018-2024, obtained from professional meteorological observation stations, ensuring its strong practical utility.
@@
To ensure a professional evaluation, we propose a more holistic evaluation framework, which assesses models across three key dimensions: i) Multi-Time Scale Prediction Evaluation: Evaluates a model’s rainfall prediction capability across different combinations of input and output sequence lengths. ii) Multi-Forecast Resolution Evaluation: It measures a model’s ability to predict rainfall at various temporal granularities. iii) Extreme Rainfall Event Evaluation: It focuses on a model’s performance in forecasting sudden, high-intensity rainfall events.
@@
Through a comprehensive evaluation, we identify that existing models often overlook the zero-inflation and temporal decay characteristics, compared to the widely acknowledged non-stationarity of time series data [10, 11, 12, 13]. To address these limitations, we design the BFPF to reinforce its sensitivity to rainfall patterns and recent temporal information. Experimental results validate the effectiveness of our approach, offering a new perspective for adapting time series models to precipitation nowcasting.
@@
  • Professional Dataset for Precipitation Nowcasting: Our benchmark is constructed from data collected at hourly intervals between 2018 and 2024 from over 140 GNSS stations globally. It covers six key variables, including PWV, and is specifically curated to support precipitation nowcasting. The dataset will be continuously updated.
@@
  • Rainfall-Centric Evaluation Strategy: We design a tailored evaluation strategy from a meteorological perspective, focusing on multi-scale forecasting, multi-temporal resolution, and extreme rainfall events.
@@
  • Novel plug-and-play Module for Precipitation Nowcasting: We introduce the Bi-Focus Precipitation Forecaster, a plug-and-play module that explicitly addresses zero inflation and temporal decay in rainfall data, achieving state-of-the-art performance in extreme rainfall forecasting.
II Related Works
II-A GNSS-based Precipitation Nowcasting
In recent years, GNSS-derived Precipitable Water Vapor (PWV) has gained considerable attention for its potential in precipitation nowcasting. Yao et al. [14] proposed a method where a sharp rise in PWV can signal impending rainfall by analyzing hourly data from the Zhejiang Continuously Operating Reference Station (CORS) network between 2014 and 2015. Unlike previous work focused primarily on statistical relationships between PWV and rainfall, Profetto et al. [15] Profetto et al. proposed a novel two-step machine learning framework that combines a Random Forest (RF) model with a Long Short-Term Memory (LSTM) neural network, which was validated using data collected between 2021 and 2023 from the GNSS meteorology station located on the roof of the LaMMA Consortium in Sesto Fiorentino, Tuscany. Liu et al. [16] proposes a novel deep learning-based model for precipitation nowcasting, which integrates GNSS-derived precipitable water vapor (PWV) data with radar observations. Lu et al. [17] proposed an enhanced precipitation nowcasting model, RSG-GAN, which integrates radar QPE, GOES-16 SWD, and GNSS ZTD data to improve forecasting accuracy over the U.S. west coast. Yin et al. [18] proposed the approach utilized machine learning algorithms to predict lightning occurrences up to 30 minutes in advance.
However, most of these studies are based on a limited number of GNSS stations in local regions, or they focus solely on establishing predictive relationships between PWV as a single variable and rainfall. To enable large-scale validation and fully leverage additional meteorological variables, it is necessary to extend these approaches beyond local datasets and single-variable models.
II-B Benchmarks for Time-Series Forecasting
Time series models play a crucial role in many fields, and a variety of benchmark datasets and evaluation frameworks have been developed to standardize performance assessment and ensure comparability across studies.
@@ II-B Benchmarks for Time-Series Forecasting
For instance, FinTSB [19] emphasizes diversity, standardization, and real-world relevance in financial forecasting. Physiome-ODE [20] introduces irregularly sampled ODE-based biological datasets for IMTS evaluation. Cherry-Picking [21] warns against dataset bias and calls for more representative evaluations. TSFM-Bench and GIFT-Eval [22] assess foundation models in zero-, few-, and full-shot regimes. TFB [23] and TSPP [24] propose unified pipelines to ensure fair and reproducible forecasting. LargeST [25] offers a long-term, large-scale traffic dataset with rich metadata to evaluate deep models in realistic settings.
Despite these advancements, the field of precipitation nowcasting still lacks a comprehensive and standardized benchmarking framework. The absence of such a framework not only hinders fair and reproducible comparisons between models but also limits the systematic evaluation of model generalization under diverse meteorological conditions, which is critical for real-world deployment and operational forecasting.
II-C Benchmark for precipitation Nowcasting
PostRainBench [4] introduced a comprehensive multi-variable numerical weather prediction (NWP) post-processing benchmark with a temporal resolution of 3 hours. However, it does not include PWV and is therefore unsuitable for nowcasting applications. RainBench [26] provides a large-scale, multi-modal benchmark using SimSat, ERA5, and IMERG for global precipitation forecasting. Rodriguez Rivero et al. Shi et al. [27] proposed both a new model and a benchmark for precipitation nowcasting based on radar echo maps from the Hong Kong Observatory. Ana et al. [6] provides a comprehensive review of deep learning-based precipitation forecasting methods that utilize multi-source observational data, such as radar reflectivity and satellite imagery. However, in the domain of precipitation nowcasting based on GNSS-PWV, a comprehensive and systematic benchmark has yet to be established.
The aforementioned related works reveal that current research in GNSS-based precipitation nowcasting is predominantly focused on localized regions, with most studies concentrating on single-factor analysis (e.g., PWV), and a notable lack of research on global-scale, multi-variable integration. Additionally, while significant advancements have been made in time series forecasting across various domains, there remains a scarcity of studies specifically addressing the application of time series models for GNSS-based precipitation nowcasting. Given the increasing use of deep learning models, particularly time series forecasting models, in the field of GNSS precipitation nowcasting, there is a pressing need to construct a globally representative, multi-variable integrated dataset and to establish a robust evaluation framework based on deep learning models to facilitate the further development and application of this area of research.
III RainfallBench
@@
RainfallBench is structured into three main components: the data layer, the model layer, and the evaluation layer. The overall framework is illustrated in Figure 1. In the following sections, Section III-A presents a formal problem definition for GNSS-based precipitation nowcasting within the context of time series forecasting. Sections III-B, III-C, and III-D describe the components of the data layer, Section III-F details the model layer, and Section F covers the evaluation layer.
III-A Problem Definition
@@
We formulate precipitation nowcasting as a multivariate-to-univariate time series prediction task. Given a sequence of historical observations comprising both meteorological variables and past rainfall values, the goal is to predict future rainfall over a fixed horizon.
@@
where $\mathbf{x}_{t}$ includes meteorological factors (e.g., temperature, humidity, wind) and the rainfall measurement at time $t$, and $D$ denotes the number of input variables and $T$ represents the length of the input sequence
@@
III-B Data Collection
RainfallBnech integrates high-frequency PWV from the Nevada Geodetic Laboratory (NGL), high-resolution auxiliary meteorological data from ERA5-Land reanalysis, and global precipitation data from the GPM IMERG Final Precipitation product. The detailed information of the three data sources is summarized in the Table I.
TABLE I: The details of RainBench
Source Product Key Parameter(s) Used Spatial Resolution Temporal Resolution
NGL Troposphere Products PWV Station-wise 5 minutes
ERA5-Land Surface Pressure, 2m Temperature, win speed 0.1° x 0.1° 1 hour
GPM IMERG Precipitation 0.1° x 0.1° 30 minutes
III-B1 Data Source
GNSS PWV. The foundational atmospheric measurements for RainfallBnech are sourced from the tropospheric products generated by the Nevada Geodetic Laboratory (NGL) at the University of Nevada. NGL stands as a world leader in the processing of raw GNSS data, providing products for a vast global network that encompasses over 19,000 stations. This unparalleled global coverage enable RainfallBnech to represent a wide array of climatological and geographical regimes. NGL employs state-of-the-art processing methodologies, utilizing the GipsyX software suite developed at NASA’s Jet Propulsion Laboratory and adhering to the latest standards and reference frames from the International GNSS Service (IGS). This ensures the highest possible quality and consistency in the derived products. For the RainfallBnech dataset, we utilize the PWV variable, which are available at high temporal resolutions, including a 5-minute sampling rate for many stations. This high frequency is essential for capturing the rapid temporal evolution of atmospheric water vapor that often precedes precipitation events.
GPM. The ground-truth precipitation data for RainfallBench is sourced from the Integrated Multi-satellitE Retrievals for GPM (IMERG) final product, specifically Version 07. The Global Precipitation Measurement (GPM) mission is an international satellite constellation designed to provide next-generation observations of rain and snow worldwide. The IMERG Final Run product is selected because it is widely regarded as the highest-quality, research-grade satellite precipitation dataset available. It use the incorporation and calibration of the satellite estimates with data from the Global Precipitation Climatology Centre’s (GPCC) network of monthly surface rain gauges. This gauge-correction step significantly reduces biases and improves the overall accuracy of the precipitation estimates, making it the most suitable choice for a benchmark dataset where the quality of the target variable is paramount. The IMERG Final Run provides quasi-global (typically 60°N-S) precipitation estimates at a high spatial resolution of 0.1° x 0.1° and a half-hourly temporal resolution. This fine spatio-temporal sampling is critical for capturing the often localized and short-lived nature of convective rainfall events, which might be missed by coarser products. The specific variable used from the product is precipitation, which provides the calibrated precipitation rate in units of mm/hr.
ERA5-land. To provide auxilliray meteorological information, surface pressure, temperature and wind speed data are required. For this purpose, RainfallBench incorporates data from ERA5-Land, a global atmospheric reanalysis product generated by the European Centre for Medium-Range Weather Forecasts (ECMWF). ERA5-Land is a replay of the land component of the flagship ERA5 reanalysis, produced using the land surface model. ERA5-Land offers several key advantages that make it the ideal choice for this application. Its primary benefit is its high spatial resolution of approximately 9 km (0.1° x 0.1°), a significant enhancement over the  31 km grid of the standard ERA5 product. This finer grid is crucial for providing more accurate estimates of surface conditions at the specific locations of the GNSS stations. Furthermore, ERA5-Land provides data at an hourly temporal resolution, which aligns well with the high-frequency nature of the GNSS observations and the need to capture diurnal cycles in atmospheric variables. The dataset also provides a long and consistent historical record, with data available from 1950 to within a few days of the present, enabling the construction of long time series dataset.
III-B2 Data Processing
A critical procedure in creating a multi-source dataset like RainfallBnech is the spatial-temporal alignment of data that exist on different spatial and temporal grids.
Temporal Alignment. All data streams are aligned to a common hourly temporal grid. The 5-minute NGL ZTD data and 30-minute GPM IMERG data are sampled to produce hourly values centered on the hour, matching the native temporal resolution of the ERA5-Land data.
Spatial Alignment. For the ERA5-Land data, which represent smooth, continuous meteorological fields, a bilinear interpolation method is used. This method estimates the value at the precise latitude and longitude of a given GNSS station by taking a distance-weighted average of the values from the four nearest ERA5-Land grid cells. This approach provides a more accurate local estimate than simply selecting the value of the single nearest grid cell.
In contrast, for the GPM IMERG precipitation data, a nearest neighbor method is employed. The rainfall value assigned to a GNSS station is the value from the GPM grid cell whose centroid is closest to the station’s coordinates. Precipitation, especially from convective storms, is a highly discontinuous and highly variational field. Using an interpolation method like the bilinear technique would artificially smooth the data, averaging out high-intensity rainfall cores. This would severely underestimate extreme precipitation events and compromise the dataset’s primary utility for nowcasting heavy rain. The nearest neighbor approach, while simpler, better preserves the magnitude and location of rainfall extreme value.
Figure 2: Global distribution of 140 selected GNSS stations from the proposed RainfallBench dataset across seven continents, ensuring balanced spatial coverage for evaluating precipitation forecasting models.
III-C Quality Control
To ensure data quality, we selected the 20 stations with the highest data completeness from each continent, each of which provides continuous records covering the full period from January 1, 2018, to January 1, 2024, spanning six complete years. For intermediate missing data, we applied a differentiated interpolation strategy to fill in the gaps. For continuous variables such as PWV, T2M, SP, wind speed, and relative humidity, we applied linear interpolation to accurately preserve their spatiotemporal continuity and evolving trends. For precipitation data, forward filling was used to effectively address its discontinuous nature, thus constructing a scientifically sound and robust high-quality dataset. The geographic distribution of the 140 stations is shown in Figure 2.
III-D Datasets Analysis
III-D1 Data Overview
Figure 3: Distribution of meteorological variables at the HKST station from 2018 to 2024. The dataset contains six variables per hourly observation.